A ball is thrown upward from ground with an initial speed of 25 m/s, at the same instant, another ball is dropped from a building 15 meters high. a.) How long will the two balls be at the same height? b.) Calculate the height where the two balls meets.
1
Expert's answer
2019-08-29T09:39:53-0400

Equations of motion of balls


"y_1=25t-5t^2\\
y_2=15-5t^2"

When two balls are at the same height


"y_1=y_2"

So


"25t-5t^2=15-5t^2"

We obtain

(a)


"t=\frac{15}{25}=0.6\:\rm{s}"

(b)


"h=y_1=y_2=15-5\times 0.6^2=13.2\:\rm{m}"